OpenStudy (piercetheveil47):

What is the volume of a sphere with a radius of 3 inches? 12π cubic inches 24π cubic inches 36π cubic inches 48π cubic inches

OpenStudy (compassionate):

\[V = \frac{ 4 }{ 3 }\pi r^3\] \[V = \frac{ 4 }{ 3 }(3.14) \times (3^3) \]

OpenStudy (piercetheveil47):

can you stay while i work it out to see if I'm correct? :)

OpenStudy (compassionate):

Sure. Just multiply 3.14 and 3^3, then multiply all that by 4/3

OpenStudy (piercetheveil47):

its C, correct??

OpenStudy (compassionate):

Yup